OpenStudy (legomyego180):

Taylor series of Cos(6x) in summation notation?

OpenStudy (legomyego180):

I keep getting alternating zeroes

OpenStudy (legomyego180):

When plugging in \[f^n(0)=\cos(6x)\] \[f(0)=1 |_{n=0}\] \[f(0)=0|_{n=1}\] \[f(0)=-6^2|_{n=2}\] \[f(0)=0|_{n=3}\] \[f(0)=6^4|_{n=4}\] Sorry if my notation is weird. Im just not sure how to write this in summation notation

OpenStudy (legomyego180):

@TheSmartOne

OpenStudy (legomyego180):

Im assuming the pattern would continue like this.

OpenStudy (legomyego180):

One of my answers has \[\sum_{k=0}^{\frac{ n }{ 2 }}\] as Sigma. I have a feeling this might take care of the alternating zeroes, but I cant tell for sure

jimthompson5910 (jim_thompson5910):

\(\Large f^{(n)}(x)\) is the nth derivative of f(x), ie, how many derivatives you apply to f(x). \[\Large f^{(0)}(x) = f(x) = \cos(6x)\] \[\Large f^{(1)}(x) = \frac{d}{dx}\left[f(x)\right] = -6\sin(6x)\] \[\Large f^{(2)}(x) = \frac{d}{dx}\left[f^{(1)}(x)\right] = -36\cos(6x)\] \[\Large f^{(3)}(x) = \frac{d}{dx}\left[f^{(2)}(x)\right] = 216\sin(6x)\] \[\Large f^{(4)}(x) = \frac{d}{dx}\left[f^{(3)}(x)\right] =1296\cos(6x)\] Hopefully you see a pattern emerging?

OpenStudy (legomyego180):

But dont I need to plug 0 in for x?

jimthompson5910 (jim_thompson5910):

yes, that will generate the coefficients (well part of them anyway)

OpenStudy (legomyego180):

so the pattern is 6^n

jimthompson5910 (jim_thompson5910):

that's part of it

OpenStudy (zarkon):

just do Cos(x) and then replace x by 6x